Radon is a naturally occurring radioactive gas found in the ground and in building materials. It is easily inhaled and emits alpha particles when it decays. Cumulative radon exposure is a significant risk factor for lung cancer.
1. Calculate the density of radon at 325 K and 0.950 atm of pressure. answer in g/L
2. Are radon concentrations likely to be greater in the basement or on the top floor of a building?
1
Expert's answer
2019-12-10T07:44:52-0500
  1. Use Ideal Gas Law to calculate density of Radon at temperature and pressure given:
"PV = nRT"

"PV=\frac{m}{M}RT"

"\frac{m}{V}= \frac{PM}{RT}"

"d = \frac{PM}{RT} = \frac{0.950 \times 222}{0.082 \times 325} = 7.91 \frac{g}{L}"

2 . As M(R) is bigger than of the air (M(air)= 29g/mol), Radon is likely to accumulate in the basement of a building.